PFB2PFA(1) PFB2PFA(1) NAME pfb2pfa−convertatype1pfbfile(binaryMSDOS)intoapfa(ASCII) SYNOPSIS pfb2pfa[-v]pfbfile[pfafile] DESCRIPTION ProgramconvertsabinaryMSDOSrepresentationforatype1PostScriptfontintoareadable ASCIIversion.TheMSDOSnewline(\r)isconvertedintotheUNIXnewline(\n).Theoutputis writteninafilewhosenameisthenamethatisprovidedonthecommandlineorthebasenameof theinputfileplusextension".pfa". Withthe-voptionyougetsomeinformationaboutwhattheprogramisdoing. AUTHOR PietTutelaers rcpt@urc.tue.nl TeX 1